# Masked Attribute Description Embedding for Cloth-Changing Person Re-identification

### v2

@@ 4.6.2 Number of Stages and TrV Blocks
In this section, we discuss the layering situation of EVA02-large, conduct experiments with MADE, and fuse attribute description features that remove clothing information with low-level to high-level features of images. We tried several stratification scenarios based on experience and summarized the experimental results on the clothes-changing setting of PRCC in Table 5. We can observe that the best results are achieved when the number of stages is three, and the number of layers is 10, 10, and 4, respectively.
@@
Table 5: Ablation studies of the different number of stages and TrV blocks for MADE in clothes-changing setting on PRCC. Bold numbers the top score.
# Stage # Layer (24) PRCC rank-1 mAP 2 [12, 12] 57.6 51.5 2 [10, 14] 48.3 45.7 3 [8, 8, 8] 63.7 57.6 3 [8, 12, 4] 64.4 60.0 3 [10, 10, 4] 67.1 62.7 4 [6, 6, 6, 6] 54.0 49.0 4 [4, 4, 8, 8] 54.5 51.7
5 Conclusion
@@
We propose the Masked Attribute Description Embedding (MADE) method, which integrates a person’s visual appearance with attribute description in CC-ReID. The modeling of volatile clothing-sensitive information, including color and type, is challenging and not conducive to identifying persons in CC-ReID. To address this, we introduce multi-modal attribute description information in CC-ReID, which is more obvious and easier to extract and edit than skeletons or contours in original images. We extract descriptions suitable for persons’ images using an attribute detection model, mask the variable clothing and color information, and embed it into the image features, compelling the model to discard clothing information. Subsequently, MADE connects and embeds the masked attribute description features encoded by Linear Projection into Transformer blocks at different levels, fusing them with low-level to high-level features of the image. By mapping different feature spaces to a shared latent space, attribute description can be fused with image features, enabling the model to capture the associated information between images and descriptions effectively. We conducted experiments on PRCC, LTCC, Celeb-reID-light, and LaST. Extensive experiments have demonstrated that MADE can effectively utilize persons’ description information to improve the performance of cloth-changing person re-identification and performs well compared to state-of-the-art methods.
@@
In the future, we intend to expliot the Large Language Models (LLM) to generate better attribute descriptions, which could help futher improve the generalization ability of our method for cloth-changing person re-identification. We will also explore the possibility of our masked attribute description strategy in other cross-modality person re-identification tasks, such as visible-infrared person Re-ID, cross-resolution person Re-ID, and sketch based person Re-ID, because there are also certain attributes which remain unchanged across the modalities.

### v3

@@ IV-E3 Number of Stages and TrV Blocks
In this section, we discuss the layering situation of EVA02-large, conduct experiments with MADE, and fuse attribute description features that remove cloth information with low-level to high-level features of images. We tried several stratification scenarios based on experience and summarized the experimental results on the cloth-changing setting of PRCC in Table VII. We can observe that the best results are achieved when the number of stages is three, and the number of layers is 10, 10, and 4, respectively.
@@
TABLE VII: Ablation studies of the different number of stages and TrV blocks for MADE in cloth-changing setting on PRCC. Bold numbers the top score.
# Stage # Layer (24) PRCC rank-1 mAP 2 [12, 12] 57.6 51.5 2 [10, 14] 48.3 45.7 3 [8, 8, 8] 63.7 57.6 3 [8, 12, 4] 64.4 60.0 3 [10, 10, 4] 66.2 63.0 4 [6, 6, 6, 6] 54.0 49.0 4 [4, 4, 8, 8] 54.5 51.7
V Conclusion
@@
We propose the Masked Attribute Description Embedding (MADE) method, which integrates a person’s visual appearance with attribute description in CC-ReID. The modeling of volatile cloth-sensitive information, including color and type, is challenging and not conducive to identifying persons in CC-ReID. To address this, we introduce multi-modal attribute description information in CC-ReID, which is more obvious and easier to extract and edit than skeletons or contours in original images. We extract descriptions suitable for personal images using an attribute detection model, mask the variable cloth and color information, and embed it into the image features, compelling the model to discard cloth information. Subsequently, MADE connects and embeds the masked attribute description features encoded by Linear Projection into Transformer blocks at different levels, fusing them with low-level to high-level features of the image. By mapping different feature spaces to a shared latent space, attribute description can be fused with image features, enabling the model to capture the associated information between images and descriptions effectively. We conducted experiments on PRCC, LTCC, Celeb-reID-light, and LaST. Extensive experiments have demonstrated that MADE can effectively utilize personal description information to improve the performance of cloth-changing person re-identification and performs well compared to state-of-the-art methods.
@@
In the future, we intend to expliot the Large Language Models (LLM) to generate better attribute descriptions, which could help futher improve the generalization ability of our method for cloth-changing person re-identification. We will also explore the possibility of our masked attribute description strategy in other cross-modality person re-identification tasks, such as visible-infrared person ReID, cross-resolution person ReID, and sketch based person ReID, because there are also certain attributes which remain unchanged across the modalities.